A cage is lowered down a pit shaft, 600m deep in 46 seconds. What is its average speed in meters per second and kilometers per hour?
1
Expert's answer
2019-12-16T11:01:29-0500
"v=\frac{600}{46}=13\frac{m}{s}"

"v=\frac{600}{46}3.6=47\frac{km}{h}"